# MoCrop: Training Free Motion Guided Cropping for Efficient Video Action Recognition

### v1

@@ 3.4 Ablation Studies
(c) Enhancing Prior Works (on UCF101, I-frames only)
CoViAR [8] Baseline (Reported in [8]) 87.7 11.6
@@
(a) PullUps
(b) ApplyEyeMakeup
(c) JumpRope
(d) Archery
(e) BalanceBeam
(f) HighJump
Fig. 2: Visualization of MoCrop (green box) vs. center cropping (red box), alongside motion-density (MD) maps. Each subfigure presents two views in order: RGB frame with boxes (left) and the MD map (right), shown across six UCF101 classes.
3.5 Visualization
Fig. 2 presents qualitative examples of adaptive cropping with MoCrop. For each class, the left panel shows an RGB frame with the green box predicted by MoCrop and the red fixed center-crop box; the right panel shows the corresponding motion-density (MD) map. Across six UCF101 classes (PullUps, ApplyEyeMakeup, JumpRope, Archery, BalanceBeam, and HighJump), MoCrop places a tighter crop around the actor and the regions with salient motion (e.g., the torso during pull-ups or the take-off area in high jump). In contrast, the center crop often includes large static background or truncates the subject (notably in Archery and BalanceBeam). The MD maps on the right concentrate energy near the actor and align with the green crop, illustrating the signal that guides our policy.
4 Discussion and Conclusion
We present MoCrop, a lightweight, training-free module using compressed-domain motion vectors to generate a single, clip-level crop of motion-dense regions. By focusing inference, MoCrop reduces resolution and FLOPs, often boosting accuracy, and integrates with existing backbones without retraining.
@@ 4 Discussion and Conclusion
Limitations. MoCrop relies on codec-provided motion vectors (e.g., H.264) and is less informative for raw videos unless motion estimation is run in parallel. This dependency is acceptable since most datasets and streaming pipelines use compressed video. Large camera motion or scene-wide dynamics can bias the density map; mitigations include global-motion suppression, temporal smoothing, and combining motion vector magnitude with residual cues.

### v2

@@ IV-D Ablation Studies
(b) Enhancing CoViAR (UCF101, I-frames only)
CoViAR Baseline (Reported in [18]) 87.7 11.6
@@
Comparison with Cropping Strategies. MoCrop significantly outperforms standard cropping techniques across multiple crop ratios (Table III). At 90% crop ratio, both Random and Center crops offer slight improvements over the baseline (80.4% for both). However, as the crop ratio decreases to 50%, these fixed strategies show severe degradation (73.5% and 71.4%). In contrast, our adaptive method achieves 81.4% accuracy with an adaptive crop ratio, demonstrating superior motion-guided region selection over static approaches across all settings.
Enhancing Prior Works. To show its versatility, we applied MoCrop to CoViAR, a well-known compressed-domain method (Table III). The results are compelling. In the efficiency setting, MoCrop improves CoViAR’s accuracy to 88.5% while reducing its FLOPs from 11.6 GFLOPs to just 8.5 GFLOPs. In the attention setting, it boosts the accuracy to an impressive 89.2%. This shows that MoCrop is not just a standalone module but can also serve as a powerful enhancer for other methods in the field.
IV-E Qualitative Analysis
Fig. 2 visualizes the complete MoCrop pipeline across six representative UCF101 action classes. Each row demonstrates how raw MVs are progressively refined through MD (percentile filtering), MCS (importance sampling), and MGS (grid aggregation and region search) to produce focused crops. The final comparison (column f) shows that MoCrop consistently identifies tighter, action-centric regions compared to fixed center crops, removing irrelevant background while preserving the actor.
(a) Raw MVs (b) MD (c) MCS (d) MGS(G) (e) MGS(S) (f) Comparison
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Archery
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
BalanceBeam
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
BaseballPitch
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PizzaTossing
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PullUps
Figure 2: Qualitative visualization of the complete MoCrop pipeline across six representative UCF101 action classes. Each row shows one video processed through six stages: (a) Raw MVs: MVs extracted from H.264 compressed video, visualized as arrows showing magnitude and direction. (b) MD (Denoised): Top-1% MVs retained after percentile-based filtering. (c) MCS (Sampled): MVs after weighted importance sampling (10% of denoised MVs, biased toward high-magnitude motion). (d) MGS - Grid: Motion-density heatmap aggregated on $16{\times}9$ spatial grid. (e) MGS - Search: Optimal region identified by dual-objective scoring (white box overlaid on heatmap). (f) Comparison: Overlay comparison of Motion-Aware Crop (green box) vs. Center Crop (red box) on the same RGB frame, demonstrating MoCrop’s ability to focus on action-relevant regions.
(a) Motion-density (b) Overlay (c) Comparison
Refer to caption Refer to caption Refer to caption
Archery: Bright-corner artifact from camera shake.
Refer to caption Refer to caption Refer to caption
BandMarching: Median rescaling biases toward near-camera region.
Refer to caption Refer to caption Refer to caption
PlayingCello: Centered actor; only bow motion detected.
Refer to caption Refer to caption Refer to caption
Rafting: Water motion fills the entire frame.
Figure 3: Failure mode analysis across five representative cases. (a) Motion-density map. (b) Heatmap overlay with search result (white box). (c) Motion-Aware Crop (green) vs. Center Crop (red). See text for detailed analysis.
V Discussion and Conclusion
We present MoCrop, a lightweight, training-free module using compressed-domain MVs to identify motion-dense regions for adaptive cropping. By focusing inference, MoCrop reduces resolution and FLOPs, often boosting accuracy, and integrates with existing backbones without retraining.
@@ V Discussion and Conclusion
Limitations. Fig. 3 illustrates five failure modes where MoCrop’s motion-density approach encounters challenges. MoCrop relies on codec-provided MVs (e.g., H.264) and is less informative for raw videos unless motion estimation is run in parallel. This dependency is acceptable since most datasets and streaming pipelines use compressed video.
Specifically, the following scenarios can bias the density map: (1) Camera shake (Archery): When the cameraman shakes the lens, spurious high-magnitude MVs appear (e.g., bright regions in corners), dominating the density map and drawing the crop away from the actual actor. (2) Perspective effects (BandMarching): When subjects move uniformly across the frame, perspective makes closer objects move faster. Since our pipeline rescales MVs by clip-wise median magnitude for robustness, this causes the near-camera region to dominate, yielding incorrect crops. (3) Stationary actors (PlayingCello): When actors are already centered with minimal body translation, only small regions (e.g., the bow) exhibit motion. The density map localizes these regions but fails to cover the full actor, offering little advantage over center cropping. (4) Scene-wide motion (Rafting): Turbulent water or dynamic backgrounds generate MVs across the entire frame, producing nearly uniform density maps. MGS then selects regions close to the full frame, providing no effective spatial focus.
Mitigations include global-motion suppression, temporal smoothing, and combining MV magnitude with residual cues. These failure modes occur in a minority of UCF101 action classes and primarily arise when motion does not reliably indicate actor location.